OpenStudy (anonymous):

Find the LCM: x^2-9 and 3x-9

OpenStudy (anonymous):

help!!!

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

3(x-3)(x+3) is divisible by both of those and there is nothing small that is divisible by both.

OpenStudy (anonymous):

'small' should have been 'smaller' in previous post

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

x² - 9=(x-3)(x+3) 3x - 9=3(x-3) so the LMC (not just least, the only one) is (x-3)

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

1) 3(x - 3) 2) 3(x + 3) 3) 3(x -3)(x + 3) 4) (x - 3)(x + 3)

OpenStudy (anonymous):

are these 1, 2, 3 the options or what ?

OpenStudy (anonymous):

options i have o choose from

OpenStudy (anonymous):

to

OpenStudy (anonymous):

Well, none of these options are the LMC of x²-9 and 3x-9

OpenStudy (anonymous):

I beg to differ. The LMC is number 3.